A consulting company employs 5 people who each work an average of 2,000 hours per day. year
(incl. normal overtime). The company's total costs are budgeted at NOK 6,000,000 (incl. Salary
to the consultants). Total target turnover is NOK 7,500,000.
a) What is the average cost per. hour consultants work on customer assignments? Take
the prerequisites you need to answer the question
b) What seems to be the price you on average think of invoicing per. hour. Justify your
solution.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
A consulting company
Workings Note
Number of employs                  5.00 A
Hours worked in a year          2,000.00 B
Total Hours worked        10,000.00 C=A*B
Answer a
Budgeted total costs 6,000,000.00 D
Total Hours worked        10,000.00 See C
Average cost per hour              600.00 E=D/C
Answer b
Target turnover 7,500,000.00 F
Total Hours worked        10,000.00 See C
Average invoicing per Hour              750.00 G=F/C